# MTTR-A: Measuring Cognitive Recovery Latency in Multi-Agent Systems

### v2

@@ Abstract
Ensuring cognitive stability in autonomous multi-agent systems (MAS) is a central challenge for large-scale, distributed AI. While existing observability tools monitor system outputs, they cannot quantify how rapidly agentic workflows recover once reasoning coherence has been lost. We adapt classical reliability metrics-Mean Time-to-Recovery (MTTR), Mean Time Between Failures (MTBF), and related ratios-into the cognitive domain, defining MTTR-A (Mean Time-to-Recovery for Agentic Systems) as a runtime measure of cognitive recovery latency. MTTR-A quantifies the time required for a MAS to detect reasoning drift and restore consistent operation, capturing the recovery of reasoning coherence rather than infrastructural repair.
A benchmark simulation using the AG News corpus and the LangGraph orchestration framework was conducted, modeling recovery latencies across multiple reflex modes. Automated reflexes restored stability within approximately 6 s on average, while human-approval interventions required about 12 s. Across 200 runs, the median simulated MTTR-A was $6.21\pm 2.14$ s, $\mathrm{MTBF}\!\approx\!6.73\pm 2.14$ s, and $\mathrm{NRR}\!\approx\!0.08$, demonstrating measurable runtime resilience across reflex strategies.
By formalizing recovery latency as a quantifiable property of distributed reasoning-and deriving reliability bounds linking recovery time and cognitive uptime-this work establishes a foundation for runtime dependability in agentic cognition, transforming cognitive recovery from an ad-hoc process into a standardized, interpretable performance metric. [^1] Code, data, and experimental configurations are available at https://github.com/BarakOr1/MTTR-A.
Keywords: multi-agent systems, cognitive reliability, agentic orchestration, LLM agents, runtime stability, fault tolerance, AI governance, reflex control, LangGraph
1. Introduction
@@
Motivation for System-Level Metricization:
@@
Central Claim Cognitive faults in multi-agent systems are inevitable; what defines reliability is not their absence but how quickly the system detects, isolates, and recovers from them.
@@
1.1. Our Approach: A Reliability Metricization Framework
@@
System Engineering Perspective Each agent acts as a repairable subsystem, and the orchestration layer composes these into a hierarchical reliability model where reflex actions represent components, agents represent subsystems, and the entire multi-agent workflow represents the system.
@@ 2. Related Work
Gap in Existing Frameworks Observability explains what happened; resilience metrics quantify how severe it was; but orchestration reliability determines what to do next-how the system should recover, replan, seek human approval, record an audit snapshot, or execute within a controlled sandbox.
@@
3. Method
This section formalizes the orchestration control method underlying the MTTR-A framework. We propose a reflex-oriented architecture for runtime recovery in MAS, where each cognitive fault triggers an automatic or human-mediated reflex action. The architecture defines five families of recovery reflexes and organizes them within a closed feedback loop linking telemetry, meta-monitoring, and execution (Figure 3, Table 1). Each recovery cycle produces measurable latency components from which MTTR-A and related dependability metrics are computed (see Section 4).
3.1. Reflex Taxonomy and Control Loop
@@
Table 1: Reflex taxonomy and corresponding recovery actions. Each category defines runtime mechanisms for detection, intervention, and safe continuation within the orchestration control loop of the MAS.
@@
The reflex families in Table 1 define complementary layers of recovery control, from automatic replanning to human-in-the-loop oversight and safety shutdown. Collectively, they ensure that cognitive drift, tool faults, and coordination errors are corrected without systemic collapse.
3.2. Runtime Policy and Coordination Dynamics
Within a MAS, recovery latency is shaped by the type of reflex invoked and also by the efficiency of policy selection and inter-agent coordination. The meta-monitor chooses tselects the fastest valid recovery action based on telemetry triggers, such as tool errors, confidence loss, or detected drift, while policy compliance. Agents coordinate through a shared global context, synchronizing memory and actions during recovery to prevent cascading inconsistencies. These decision and communication latencies collectively define the measured MTTR-A: the time interval from fault detection to complete restoration of coherent operation across the MAS.
@@ 3.2. Runtime Policy and Coordination Dynamics
4. Experiment and Evaluation Metrics
@@
4.1. Evaluation Metrics and Cognitive Dependability Context
@@
4.2. Formal Reliability Metricization Framework
@@
Key Metric Definitions The framework defines three core reliability metrics: MTTR-A measures how quickly reasoning recovers after drift, MTBF captures the average stable interval between faults, and NRR combines both into a single indicator of runtime reliability.
@@
Figure 6: Illustrative decomposition of MTTR-A. Conceptual example showing the relative contributions of detection ($T_{\mathrm{detect}}$), decision ($T_{\mathrm{decide}}$), and execution ($T_{\mathrm{execute}}$) latencies as defined in Equation (5). This figure illustrates the theoretical structure of recovery latency.
@@
$\pi_{\mathrm{up},sys}=\frac{\mathrm{MTBF}_{sys}}{\mathrm{MTBF}_{sys}+\mathrm{MTTR}_{sys}}=\frac{1}{1+\lambda_{sys}\mu_{sys}}\geq 1-\lambda_{sys}\mu_{sys}=\mathrm{NRR}_{sys},$ (10)
@@ 4.3. Experimental Setup
4.4. Algorithmic Procedure for Measuring MTTR-A
@@ 5. Results and Discussion
These metrics indicate that the agentic workflow maintained stable recovery cycles with roughly 10% of runtime spent on corrective actions and no simulated evidence of cascading instability. Per-mode results are summarized in Table 2.
Table 2: Experimental results on the AG News dataset (N = 200 runs). Each value represents the median recovery time per reflex mode.
@@
5.1. Latency Composition by Reflex Type
Figure 11 decomposes MTTR-A into three additive components: decomposes MTTR-A into three additive components: detection, decision, and execution latency. Execution time dominates recovery cost across all reflexes, especially in human-approval cases. The low decision latency confirms that policy selection overhead within LangGraph is negligible compared to the actual reflex execution duration.
@@ 5.1. Latency Composition by Reflex Type
MTTR-A can serve as a building block for higher-level aggregate indicators-such as a Mean Cognitive Uptime (MCU) or a Cognitive Reliability Index (CRI)-which define measurable targets for adaptive AI performance and governance.
@@
6. Conclusion and Future Work
@@
Future work will extend this foundation to real-world heterogeneous teams, integrating semantic performance metrics with runtime dependability measures to assess trade-offs between reasoning quality and recovery efficiency. Further directions include exploring governance-aware reliability objectives, safety guarantees for reflexive control loops, and alignment with emerging standards for operational AI reliability.
Appendix
@@ Appendix A Proofs
$\pi_{\mathrm{up},sys}=\frac{\mathbb{E}[U]}{\mathbb{E}[U]+\mathbb{E}[D]}=\frac{\mathrm{MTBF}_{sys}}{\mathrm{MTBF}_{sys}+\mathrm{MTTR}_{sys}}=\frac{1}{1+\lambda_{sys}\mu_{sys}}.$ (A.1)
@@ Appendix B Query Pool
  "LangGraph recovery reflexes", "agent orchestration reliability",
  "rollback sandbox audit snapshots", "tool retries and backoff",
  "consensus voting disagreement", "policy thresholds and approvals",
  "governance and risk tiers", "observability telemetry signals",
  "drift detection and confidence", "mttra and mtbf calculation",
  "normalized reliability index", "incident playbooks escalation",
  "global memory reconciliation", "safe mode fallback routes",
  "retrieval reranking grounding", "planning decomposition tools"
Each query was submitted through the reasoning-drift-recovery cycle defined in Section 4, ensuring consistent cognitive fault induction and enabling reproducible measurement of recovery latency and MTTR-A dynamics.

### v3

@@ Abstract
Ensuring cognitive stability in autonomous multi-agent systems (MAS) is a central challenge for large-scale, distributed AI. While existing observability tools monitor system outputs, they cannot quantify how rapidly agentic workflows recover once reasoning coherence has been lost. We adapt classical reliability metrics-Mean Time-to-Recovery (MTTR), Mean Time Between Failures (MTBF), and related ratios-into the cognitive domain, defining MTTR-A (Mean Time-to-Recovery for Agentic Systems) as a runtime measure of cognitive recovery latency. MTTR-A captures the time required for a MAS to detect reasoning drift and restore coherent operation, focusing on recovery of cognition rather than infrastructural repair.
A benchmark simulation using the AG News corpus and the LangGraph orchestration framework evaluates recovery latencies across reflex modes. Automated reflexes restore stability within approximately six seconds, while human-approval interventions require roughly twelve seconds. Across 200 runs, the median MTTR-A was $6.21\pm 2.14$ seconds, with MTBF $\approx 6.73\pm 2.14$ seconds and a normalized recovery ratio (NRR) of approximately 0.08.
By formalizing recovery latency as a measurable runtime property and deriving reliability bounds linking recovery time to cognitive uptime, this work establishes a foundation for runtime dependability in agentic AI systems. This work does not propose a new agent architecture or learning algorithm; instead, it focuses on system-level metricization and runtime dependability of existing agentic workflows.
Index Terms:
Multi-agent systems, cognitive reliability, agentic orchestration, runtime stability, fault tolerance, AI governance.
I Introduction
@@
Motivation for System Level Metricization:
@@
I-A Our Approach: A Reliability Metricization Framework
@@ II Related Work
III Method
This section formalizes the orchestration control method underlying the MTTR-A framework. We propose a reflex-oriented architecture for runtime recovery in MAS, where each cognitive fault triggers an automatic or human-mediated reflex action. The architecture defines five families of recovery reflexes and organizes them within a closed feedback loop linking telemetry, meta-monitoring, and execution (Figure 3, Table I). Each recovery cycle produces measurable latency components from which MTTR-A and related dependability metrics are computed (see Section IV).
III-A Reflex Taxonomy and Control Loop
@@
TABLE I: Reflex taxonomy and corresponding recovery actions. Each category defines runtime mechanisms for detection, intervention, and safe continuation within the orchestration control loop of the MAS.
@@
The reflex families in Table I define complementary layers of recovery control, from automatic replanning to human-in-the-loop oversight and safety shutdown. Collectively, they ensure that cognitive drift, tool faults, and coordination errors are corrected without systemic collapse.
III-B Runtime Policy and Coordination Dynamics
Within a MAS, recovery latency is shaped by the type of reflex invoked and also by the efficiency of policy selection and inter-agent coordination. The meta-monitor chooses the fastest valid recovery action based on telemetry triggers, such as tool errors, confidence loss, or detected drift, while policy compliance. Agents coordinate through a shared global context, synchronizing memory and actions during recovery to prevent cascading inconsistencies. These decision and communication latencies collectively define the measured MTTR-A: the time interval from fault detection to complete restoration of coherent operation across the MAS.
@@ III-B Runtime Policy and Coordination Dynamics
IV Experiment and Evaluation Metrics
@@
IV-A Evaluation Metrics and Cognitive Dependability Context
@@
IV-B Formal Reliability Metricization Framework
@@
Figure 6: Decomposition of MTTR-A. Conceptual example showing the relative contributions of detection ($T_{\mathrm{detect}}$), decision ($T_{\mathrm{decide}}$), and execution ($T_{\mathrm{execute}}$) latencies as defined in Equation (5). This figure illustrates the theoretical structure of recovery latency.
@@
$\pi_{\mathrm{up},sys}$ $=\frac{\mathrm{MTBF}_{sys}}{\mathrm{MTBF}_{sys}+\mathrm{MTTR}_{sys}}=\frac{1}{1+\lambda_{sys}\mu_{sys}}$ (10)
$\geq 1-\lambda_{sys}\mu_{sys}=\mathrm{NRR}_{sys}.$
@@ IV-C Experimental Setup
IV-D Algorithmic Procedure for Measuring MTTR-A
@@ V Results and Discussion
These metrics indicate that the agentic workflow maintained stable recovery cycles with roughly 10% of runtime spent on corrective actions and no simulated evidence of cascading instability. Per-mode results are summarized in Table II.
TABLE II: Experimental results on the AG News dataset (N = 200 runs). Each value represents the median recovery time per reflex mode.
@@
V-A Latency Composition by Reflex Type
Figure 11 decomposes MTTR-A into three additive components: detection, decision, and execution latency. Execution time dominates recovery cost across all reflexes, especially in human-approval cases. The low decision latency confirms that policy selection overhead within LangGraph is negligible compared to the actual reflex execution duration.
@@ V-A Latency Composition by Reflex Type
MTTR-A can serve as a building block for higher-level aggregate indicators, such as a Mean Cognitive Uptime (MCU) or a Cognitive Reliability Index (CRI)-which define measurable targets for adaptive AI performance and governance.
@@
VI Conclusion and Future Work
@@
Future work will extend this foundation to real-world heterogeneous teams, integrating semantic performance metrics with runtime dependability measures to assess trade-offs between reasoning quality and recovery efficiency. Further directions include exploring governance-aware reliability objectives, safety guarantees for reflexive control loops, and alignment with emerging standards for operational AI reliability. In practice, MTTR-A enables organizations to define service-level objectives (SLOs) for agentic systems based on recovery behavior rather than task accuracy alone.
@@ Appendix A Proofs
$\pi_{\mathrm{up},sys}$ $=\frac{\mathbb{E}[U]}{\mathbb{E}[U]+\mathbb{E}[D]}=\frac{\mathrm{MTBF}_{sys}}{\mathrm{MTBF}_{sys}+\mathrm{MTTR}_{sys}}$ (A.1)
$=\frac{1}{1+\lambda_{sys}\mu_{sys}}.$
@@ Appendix B Query Pool
"LangGraph recovery reflexes",
"agent orchestration reliability",
"rollback sandbox audit snapshots",
"tool retries and backoff",
"consensus voting disagreement",
"policy thresholds and approvals",
"governance and risk tiers",
"observability telemetry signals",
"drift detection and confidence",
"mttra and mtbf calculation",
"normalized reliability index",
"incident playbooks escalation",
"global memory reconciliation",
"safe mode fallback routes",
"retrieval reranking grounding",
"planning decomposition tools"
Each query was submitted through the reasoning-drift-recovery cycle defined in Section IV, ensuring consistent cognitive fault induction and enabling reproducible measurement of recovery latency and MTTR-A dynamics.